OpenStudy (anonymous):

solve the equation, cos^2(x)-cos(x)=2 in the domain of 0 is less than or equal to x and x is less than or equal to 2π.

OpenStudy (anonymous):

X=pi

OpenStudy (anonymous):

how do you get that answer?

OpenStudy (anonymous):

Cos^2(pi)=1 and cos(pi)=-1.1--1=1+1=2. I actually just did it mentally.

OpenStudy (anonymous):

Thank you so much!!!

OpenStudy (anonymous):

can you also help me with one more problem?

OpenStudy (anonymous):

Certainly!

OpenStudy (anonymous):

Great! Calculate the distance between the two points (t, u-2) and (t-2, u). Please use radical form in your solution.

OpenStudy (anonymous):

Do you know how to solve it?

OpenStudy (anonymous):

The solution is 2(2^(1/2)). We use the formula d=((x_2-x_1)^2+(y_2-y_1)^2)^(1/2) with x_2=t, x_1=t-2, y_2=u-2, and y_1=u. Sorry it took a while to reply. I'm typing on an iPad.

OpenStudy (anonymous):

No problem! Thank you so much!

OpenStudy (anonymous):

what does the underscore stand for?

OpenStudy (anonymous):

That's to indicate a subscript since I can't type that on an iPad.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so the answer is 2(2^1/2)?

OpenStudy (anonymous):

Yes. That x^(1/2) is the same as the square root of x.

OpenStudy (anonymous):

wait so my answer is 2(2√x) ???